OpenStudy (anonymous):

how to convert 1.50 micrometers^2 to m^2

OpenStudy (anonymous):

i know that 1 micometer =10^-6

OpenStudy (unklerhaukus):

well one micro meter is 1x10^-6 meters so 1 square micro meter is 1x10^-12 m^2 hence \[1.50\mu m^2=1.50\mu m^2 \times { 10^{-12} }{ m^2\over\mu m^2}\]\[=1.50 \times 10^{-12} m^2\]

OpenStudy (anonymous):

i thought that we have to do something like this: 1.50micrometerx m^2/10^-6 micrometers ^2

OpenStudy (unklerhaukus):

\[1={m \over 10^{6} \mu m}=10^{-6}{m \over \mu m}=10^{-12}{m^2 \over \mu m^2}\]

OpenStudy (anonymous):